Commissioning Notes from 20-21 May 2008 MST

Observers: D.Thompson at the LBT and A.Rakich, D.Miller in Tucson. OSA: David G.-H.

Summary

A night plagued by high wind, apalling seeing and software glitches, got some data but very little. The little field data we did get is suggesting a reasonably significant focal plane tilt. One good result was sorting out rotation of guide stars into Zemax field coordinates. Hoping for better luck tomorrow night.

Nighttime Activities

02:40 We had opened earlier, but have to close because the wind is too high. Note that there are 2 fires on the mountain at low elevation. We need to keep an eye on wind direction and close if the smoke comes over the dome.

04:25 Back open, pointto M5_0531 ptmodify to set IE = -38, CA = -89 to being star on center. Lots of coma.

04:40 pointto BS9152 GCS is failing to start guiding. We tried stopping and restarting several subsystems and it GCS is still failing consistently.

05:11 Called Tony Edgin...

05:15 Doing full TCS shutdown and restart...05:45 seems to be back up

05:49 We are now way out of focus, and not getting rotator polys.

05:55 Called D.Ashby

06:01 OK, we are back after several restarts of PCS

06:05 Problems continue, called Tony again. He reccommends calling Tom

06:10 Called Tom,

Center2 script updated, needs testing.

06:33 We were hitting the wrap limit of the rotator. This is opposite last month's behavior because of the rotation to put PA=0 on DS9 rather than on the IRTC GUI. Stay to the SE or SW to remain away from rotator limits (at PA=0).

M5_0699 to recenter: IE -62.00000 CA -82.00000

BS9177 (near M5_0699) - lots of coma, but we are on source.

07:11 Starting to work on WFS

7:30 Took images WFSC000002-11 in /Repository/AGW_Data/080521w, closed loop on-axis on BS9177

7:31 Gone to off-axis star #54 set exp to 2000 too faint, going to another star #57, didn't work out, gone to # 14

7:48 Did a quick test of "164 - Rotang" to project the y axis on the AGW onto the yaxis on M1. Did a +Rx = 5" rotation on M1, saw the star move up and at the right angle, going back to guiding

8:01 Taking 5 sets of images on star 14 after closed loop ACtO drove aberrations under 50 nm per term.

8:10 AGW camera hanged. Need to restart AGW

8:21 Taking multiple images (5) of star 14 after collimation, WFSC000021-25, stopped gone to star 19

8:32 closed due to wind

9:28 open, going back to BS9177 on-axis.

9:35 Collimating WFSC 000026-36, seeing 3"

9:54 taking series of 5 exposures on-axis wfsc000037-41

9:58 going to star 63, taking series of 5 images without collimation, R magnitude 10.96, difficulty finding star, got it

10:07 start taking 5 images WFSC 000047-51 pupil badly vignetted but will measure anyway, dome closed halfway through due to high wind

-- DavidThompson - 21 May 2008
Topic revision: r6 - 21 May 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback